# Milliarcsecond astrometric oscillations in AGN as a precursor of multi-messaging gravitational wave events

### v1

@@ 4 Toy model of SMBHB evolution
The first phase of SMBHB evolution is considered in a number of works (e.g. Kou et al., 2024), where formation of supermassive black hole binaries is considered as a result of collision of galaxies, in the nucleus of which there are single supermassive black holes. Based on the available statistical observational data, it is concluded that such binary systems are insufficient to explain the GW background detected using pulsar timing arrays (see, e.g., Agazie et al., 2023, and references therein). If we remain within the hypothesis that this GW background is generated by SMBHBs, we should assume that not all binary black hole systems are formed as a result of the merger of two galaxies. It is possible that only one of the two black holes with the mass is surrounded by a dense medium in the central region of a galaxy, whereas the second, less massive black hole of the mass (it could be, for example, a primordial black hole) is captured by a galaxy with the central BH of the mass . Such events could be more frequent than the collisions of galaxies and could generate the GW background mentioned above. SMBHBs still remain a rather elusive class of extragalactic objects not only in the observational perspective as mentioned above, but also from the theoretical standpoint: presently, their abundance and mass function are very uncertain.
@@
For the sake of simplicity, we assume in our toy model that the mass ratio of black holes in a binary system, is small. Recently, Kuznetsov (2024) (for the object OJ 287) and Titarchuk & Seifina (2024) (for the object SDSS J075217.84+193542.2) demonstrated that X-ray observations support this assumption and give , also consistent with the estimate for the source J10487143 by Kun et al., 2024 (see Section 1).
@@
Such the assumption makes it possible to describe the motion of a less massive black hole as the motion of a test body in a given field of a central more massive black hole. Then will be defined by Kepler’s Third Law:
@@
where is the gravitational constant, , , observed period, and the redshift of the object of interest. For , and redshift of the source J02041415, , we come to the following estimate of the separation between the binary system components:
@@
Such the value of separation corresponds to the angular size of the order of 1 as. This value is very close to the estimate obtained for the suggested SMBHB system in the object PKS 2131021: the separation between the components pc with the rest-frame orbital period of yr (O’Neill et al., 2022). However, this latter result is obtained by the method different from ours, namely, by radio flux density monitoring over more than 45 years. Recently, this result for PKS 2131021 has been strengthened by a discovery with the same technique of the second similar object, PKS J08050111 with the rest frame period 1.4220.005 yr (de la Parra et al., 2024).
@@
From Eq. 14, we can see that for the object J21026015, is considerably larger than . This suggests the following simplification for our toy model: the relativistic corrections to all parameters of the secondary black hole orbit are not relevant for order of magnitude estimates. In particular, this means that we can consider the central black hole to be Schwarzschild rather than the Kerr one, in other words, we can neglect, along with all other relativistic effects, those related to the rotation of the central SMBH.
@@
The next hardening phase when the less massive black hole gradually gets closer to the central black hole, can naturally be divided into two distinct stages. During the first stage (considered in subsection 4.2 below), the energy and angular momentum losses of already formed SMBHB due to the radiation of gravitational waves are negligible and the characteristic evolution time is determined by the interaction of the second, less massive black hole with matter around the more massive central black hole.
During the second stage (considered in subsection 4.3), when the two black holes are close enough to each other, the main losses of energy and angular momentum are due to with the radiation of gravitational waves.
4.2 Dynamical friction followed by interaction with accretion disk
@@
At the beginning of the first stage, the most significant role in the evolution of SMBHB is played by dynamical friction, i.e. gravitational interaction with stars in the cluster around the central SMBH. As was shown in Polnarev & Rees (1994), dynamical friction results in loss of energy and angular momentum by the secondary black hole in such a way that, in the most probable cases, the orbital eccentricity of the secondary black hole is rather small. For this reason (rather than for the reasons of simplicity only), we assume that the eccentricity is zero, i.e. the orbit of the secondary black hole is assumed to be circular.
@@
This simplification means, in particular, that the astrometric oscillations noted above are not related to the Lidov–Kozai effect in celestial mechanics (Lidov, 1962; Kozai, 1962). The application of this effect to SMBHBs deserves a separate investigation (see, e.g., Naoz, 2016; Ivanov et al., 2005) and goes beyond the scope of this work.
@@
The main problem associated with dynamical friction can be formulated as follows: will the second black hole make it in cosmological time (which is of the order yr, where is the Hubble constant) to get so close to the central black hole that dynamical friction ceases to be effective, but other factors come into play that lead to a faster convergence of black holes? The ways of observational and theoretical solutions to this problem and its relation to the problem of the last parsecs are discussed in more detail in the review of D’Orazio & Charisi (2023). We can make the following assumption, based on the conclusions made in this review: for a wide range of parameters describing black holes and star clusters around the central black hole, the answer to this question is positive, i.e. we assume (without proof) that the timescale of reproaching of the secondary black hole to the central black hole due to dynamical friction in the cluster is considerably shorter than cosmological time. In other words, we assume that the the problem of the last parsec is somehow solved. A more detailed analysis of the issue is beyond the scope of the present work.
@@
As a less massive black hole approaches a more massive one, the role of the dynamic friction considered above in the evolution of the orbit gradually decreases. Sooner or later, the interaction of a less massive black hole with the accretion disk around the central black hole begins to play a major role in this evolution. A detailed analysis of this interaction is outside of the scope of the present work and must take into account that the accretion disk around the primary SMBH is not necessarily thin, it may even be strongly twisted, and the orbit of the secondary black hole may be non-circular and arbitrarily oriented relative to the disk (see, e.g., Ivanov & Zhuravlev, 2024).
@@
However, it is possible to give a rough but sufficient for our toy model estimate for orbital evolution timescale of this interaction, (see Ivanov et al., 1999, 2015). If is the mass of the smaller black hole, and is the accretion rate onto the more massive black hole then
where and (see Genzel et al., 2024 for motivation of such the normalisation of ; see also Purohit et al., 2024; D’Orazio & Charisi, 2023).
@@
According to Eq. 15, for reasonable choice of parameters, the timescale of evolution of SMBHB due to its interaction with the accretion disk is also less than the age of the Universe, just as the timescale of dynamical friction discussed above.
@@ 4.3 Gravitational wave emission and orbital evolution of SMBHB
In this subsection, the characteristic timescale of evolution of the SMBHB orbit due to the emission of gravitational waves is compared to the corresponding characteristic timescale of orbit evolution defined by interaction with the accretion disk (see Eq. 15).
For a circular orbit and small (see, e.g., Landau & Lifshitz, 1975, Chapter 13, Section 110), the separation between two components in a binary system changes with time according to
where is the power carried away by gravitational waves:
@@
This variation of corresponds to the timescale
From , and taking into account Eq. 15 and 19, we find that the radiation of gravitational waves plays the main role in the orbital evolution of SMBHB when
As follows from Eq. 14 and 20, we cannot exclude that evolution of the source J02041415 is already determined by the radiation of gravitational waves because . The corresponding timescale of orbit evolution, as follows from (14) and (19), is:
@@
Since the statistical distributions of major physical parameters (masses and mass ratios, spins, all orbital parameters) of the SMBHBs as a class of objects largely unknown, we refrain from making any other predictions in this paper about the GW manifestations of SMBHBs for shorter time intervals. Nevertheless, in the next two subsections, we present some considerations on the possibility of detecting gravitational waves from individual SMBHBs by laser interferometer antennas like LISA (Colpi et al., 2023) (in subsection 4.4) and possible contribution of SMBHBs into the GW background detected by pulsar timing arrays (Agazie et al., 2023, and references therein) (in subsection 4.5). These considerations outline the contours of future studies of SMBHBs as multi-messenger objects.
4.4 Detection of gravitational waves generated by individual SMBHB
In this subsection, we touch briefly future detection of GWs by LISA-type detectors and give some rough constraints of SMBHB parameters for which such detection is possible.
Let us assume that a prospective GW detector can detect all gravitational waves if (i) their frequencies are in the range
and (ii) their amplitudes (such amplitude is called the strain of GW) are larger than some value (called the sensitivity of the detector).
Such an assumption is a rough replacement for the non-monotonic dependence of the sensitivity of any real gravitational wave detector on the frequency of the receiving signal – see, for example, Colpi et al. (2023) for LISA and Agazie et al. (2023) for NANOgrav. The main purpose of such the assumption is not only to simplify the restrictions given below on the parameters of SMBHBs, but also to make the physical meaning of these restrictions more clear.
@@ 4.4 Detection of gravitational waves generated by individual SMBHB
Another constraint comes from a comparison of the amplitude of the emitted GW with . To make an order of magnitude estimate, we can use the quadrupole formula (see Landau & Lifshitz, 1975, Chapter 13, Section 110). Omitting tensor indices we get
where is the second time derivative of the quadrupole moment, is the orbital angular velocity, and is the luminosity distance from SMBHB to the detector of GWs. Then one can obtain the following inequality:
@@
We have the last very simple constraint on from the well known fact that the radius of the last stable circular orbit around a Schwarzschild black is . Hence
@@
Finally, we have the following constraints on parameters of SMBHBs generating detectable GWs (see also Ivanov et al., 2015):
@@ 4.5 SMBHBs and GW background detected by pulsar timing arrays
This subsection is devoted to the assumed population of SMBHBs which contribute to the GW background radiation detected by pulsar timing arrays (Agazie et al., 2023, and references therein). Let’s assume that the population of SMBHBs is evolving with the characteristic time given by Eq. 15, and consists of objects with approximately the same parameters. The number of SMBHBs, , that could make a significant contribution to the GW background detected by pulsar timing arrays above a given sensitivity threshold and in a given frequency range (Eq. 22) is , where . As follows from Eq. 15 and 19,
The SMBHB separation still satisfies inequalities 26 and 30, while inequality 29 should be significantly modified. Indeed, for background radiation, inequality 28 should be rewritten as
@@
The constraint on can be obtained now as a condition for the compatibility of inequalities 37 and 26:
hence
@@
It is notable that parameter in our extremely simplified toy model does not appear in the above constraint for detectability of GW by pulsar timing arrays, contrary to the constraint in Eq. 33 for detectability of individual SMBHB GW generator. The characteristic time of the existence of close enough SMBHBs is proportional to (see Eq. 15), while the characteristic time when individual SMBHBs contribute to the background radiation is inversely proportional to (see Eq. 19). Hence, when we deal with the GW background, the fraction of contributing SMBHBs is proportional to (see Eq. 35). For this reason, the corresponding requirement for the amplitude of the gravitational wave, which itself is proportional to (see Eq. 29) is weakened by a factor of (see Eq. 35 and 36). The fact that the intensity of background GW radiation generated by SMBHBs does not depend on (or depends weakly in a model more accurate than the toy model presented in this paper) favours formation of SMBHBs from less numerous but more massive SMBHs and more numerous but less massive SMBHs.
@@
5 Outlook for future direct imaging studies of SMBHBs as multi-messangers
@@
As shown above, getting a decisive information on the parameters of an SMBHB as a generator of GW requires its measurable parameters obtained at the microarcsecond scale, at least indirectly. Arguably, the most convincing evidence and in-depth study opportunity would be provided by direct imaging observations of potential SMBHB objects with a microarcsecond resolution, in which the duality of the source’s morphology would be visible directly, like in several cases mentioned in Section 1 but at much smaller separations of milliparsecs – see the estimates presented in Section 4. This can be illustrated following a very convenient visual representation presented in Fig. 1 by D’Orazio & Charisi (2023): the dual AGN currently occupying the very top edge of the plot with the rest-frame periods of the order of yr should be ‘pushed down’ by five orders of magnitude to the orbital periods of the order of years in the area of VLBI astrometry.
@@
The microarcecond angular resolution is beyond reach of any existing astronomical imaging instrument or facility in any domain of the electromagnetic spectrum. The currently record-holding Earth-based VLBI instruments cannot provide such the resolution due to fundamental limitations of the available longest baselines which cannot exceed the Earth’s diameter, and the Earth’s atmosphere opacity at wavelengths shorter than about 0.7 mm. The only plausible solution for achieving the required angular resolution is placing a VLBI system in space, above the Earth’s atmosphere and with no limitations on the baseline length. Several concepts of such systems have been developed over the past decade. One of them, THEZA (TeraHertz Exploration and Zooming-in for Astrophysics) has been presented in response to the European Space Agency’s call for inputs into the prospective strategic program Voyage 2050 (Gurvits et al., 2021, 2022). It suggests a multi-element space-borne VLBI configuration able to operate at mm and sub-mm wavelengths with the angular resolution of the order of 1 as. Another concept called BHEX (Black Hole Explorer) considers an Earth–Space VLBI configuration operating at mm/sub-mm wavelengths (Johnson et al., 2024). With a bit of luck, and accepting the prediction by D’Orazio & Charisi (2023, section 3.1), the currently ongoing analysis of the RadioAstron non-imaging AGN Survey (Kovalev et al., 2020), reaching in the best cases the angular resolution of might offer one to several new SMBHB candidates. However, they will still require decisive imaging observations with future space-borne VLBI systems.
@@
Another potentially efficient multi-messaging approach to investigating the contribution of merges of black hole binaries into the GW background at frequencies higher than those explored by PTA, thus accessible to LISA, is suggested by Ellis et al. (2024). Their model reconciles the pre-JWST and JWST observations of the co-growth of masses of stellar population of galaxies and their central black holes, including SMBHBs. The quantitative estimates of GW generation by evolving SMBHBs presented in section 4 might become useful in assessing the observational perspectives of LISA and other gravitational wave instruments covering specific frequency ranges.
@@
While discussing prospects of future detections of SMBHBs as ‘visually’ binary AGN, we need to underline the difference between traditional morphological VLBI studies of jets in AGN, especially their innermost bases, and the SMBHs themselves. The latter would become accessible for direct imaging observations if the technique demonstrated by the mm-wavelength VLBI Event Horizon Telescope (EHT) in the detection of the black hole shadows in M 87∗ (EHT, 2019) and Sgr A∗ (EHT, 2022) improves its angular resolution at least by an order of magnitude. Again, this would become possible only with prospective space-borne extensions of EHT like BHEX and THEZA.

### v2

@@ 4.1 Formation and evolution of SMBHBs
The first phase of SMBHB evolution is considered in a number of works (e.g. Kou et al. 2024), in which formation of SMBHBs is considered to be a result of collision of galaxies, in the nucleus of which there are single SMBHs. Based on the available statistical observational data, it is concluded that such binary systems are insufficient to explain the GW background detected using PTAs (see, e.g. Agazie et al. 2023, and references therein). If we remain within the hypothesis that this GW background is generated by SMBHBs, we should assume that not all binary black hole systems are formed as a result of the merger of two galaxies. It is possible that only one of the two black holes, with the mass , is surrounded by a dense medium in the central region of a galaxy, whereas the second, less massive black hole with the mass (it could be, for example, a PBH) is captured by a galaxy with the central BH of the mass . Such events could be more frequent than collisions of galaxies and could generate the GW background mentioned above. SMBHBs still remain a rather elusive class of extragalactic objects not only from the observational perspective mentioned above, but also from the theoretical standpoint: presently, their abundance and mass function are very uncertain.
@@
For the sake of simplicity, we assume in our toy model that the mass ratio of black holes in a binary system, , is small. Recently, Kuznetsov (2024) (for the object OJ 287) and Titarchuk & Seifina (2024) (for the object SDSS J075217.84+193542.2) demonstrated that X-ray observations support this assumption and give , which is also consistent with the estimate for the source J10487143 by Kun et al. 2024 (see Section 1).
@@
Such an assumption makes it possible to describe the motion of a less massive black hole as the motion of a test body in a given field of a more massive central black hole. Then will be defined by Kepler’s third law:
@@
where is the gravitational constant, , , the observed period, and the redshift of the object of interest. For , , and the redshift of the source J02041415, , we come to the following estimate of the separation between the binary system components:
@@
Such a value of separation, , corresponds to an angular size of the order of 1 as. This value is very close to the estimate obtained for the suggested SMBHB system in the object PKS 2131021: the separation between the components pc with the rest-frame orbital period of yr (O’Neill et al. 2022). However, this latter result is obtained by a method different from ours; namely, by radio flux density monitoring over more than 45 years. Recently, this result for PKS 2131021 has been strengthened by a discovery with the same technique of a second, similar object, PKS J08050111, with the rest frame period 1.4220.005 yr (de la Parra et al. 2024).
@@
From Eq. 14, we can see that for the object J21026015, is considerably larger than . This suggests the following simplification for our toy model: the relativistic corrections to all parameters of the secondary black hole orbit are not relevant for order-of-magnitude estimates. In particular, this means that we can consider the central black hole to be Schwarzschild rather than the Kerr one. In other words, we can neglect, along with all other relativistic effects, those related to the rotation of the central SMBH.
@@
The next hardening phase, when the less massive black hole gradually gets closer to the central black hole, can naturally be divided into two distinct stages. During the first stage (considered in subsection 4.2 below), the energy and angular momentum losses of already formed SMBHB due to the radiation of GWs are negligible and the characteristic evolution time is determined by the interaction of the second, less massive black hole with matter around the more massive central black hole. During the second stage (considered in subsection 4.3), when the two black holes are close enough to each other, the main losses of energy and angular momentum are due to the radiation of GWs.
4.2 Dynamical friction followed by interaction with the accretion disc
@@
At the beginning of the first stage, the most significant role in the evolution of the SMBHB is played by dynamical friction, i.e. gravitational interaction with stars in the cluster around the central SMBH. As was shown in Polnarev & Rees (1994), dynamical friction results in loss of energy and angular momentum by the secondary black hole in such a way that, in the most probable cases, the orbital eccentricity of the secondary black hole is rather small. For this reason (rather than only for simplicity), we assume that the eccentricity is zero, i.e. the orbit of the secondary black hole is assumed to be circular.
@@
This simplification means, in particular, that the astrometric oscillations noted above are not related to the Lidov–Kozai effect in celestial mechanics (Lidov 1962; Kozai 1962). The application of this effect to SMBHBs deserves a separate investigation (see, e.g. Naoz 2016; Ivanov et al. 2005) and goes beyond the scope of this work.
@@
The main problem associated with dynamical friction can be formulated as follows: whether the second black hole will manage in cosmological time (which is of the order yr, where is the Hubble constant) to get so close to the central black hole that dynamical friction ceases to be effective, but other factors come into play that lead to a faster convergence of black holes. Observational and theoretical solutions to this problem and its relation to the problem of the last parsecs are discussed in more detail in the review of D’Orazio et al. (2024). We can make the following assumption, based on the conclusions made in this review: for a wide range of parameters describing black holes and star clusters around the central black hole, the answer to this question is positive, i.e. we assume (without proof) that the timescale of reproaching of the secondary black hole to the central black hole due to dynamical friction in the cluster is considerably shorter than cosmological time. In other words, we assume that the problem of the last parsec is somehow solved. A more detailed analysis of the issue is beyond the scope of the present work.
@@
As a less massive black hole approaches a more massive one, the role of the dynamic friction considered above in the evolution of the orbit gradually decreases. Sooner or later, the interaction of a less massive black hole with the accretion disc around the central black hole begins to play a major role in this evolution. A detailed analysis of this interaction is outside of the scope of the present work and must take into account that the accretion disc around the primary SMBH is not necessarily thin, and may even be strongly twisted, and that the orbit of the secondary black hole may be non-circular and arbitrarily oriented relative to the disc (see, e.g. Ivanov & Zhuravlev 2024).
@@
However, for our toy model it is possible to give a rough but sufficient estimate of the orbital evolution timescale of this interaction, (see Ivanov et al. 1999, 2015). If is the mass of the smaller black hole, and is the accretion rate onto the more massive black hole, then
@@
where and (see Genzel et al. 2024 for the motivation of such a normalisation of ; see also Purohit et al. 2024; D’Orazio et al. 2024). According to Eq. 15, for a reasonable choice of parameters, the timescale of the evolution of a SMBHB due to its interaction with the accretion disc is also less than the age of the Universe, just as the timescale of dynamical friction discussed above is.
@@ 4.3 Gravitational wave emission and orbital evolution of SMBHB
In this subsection, the characteristic timescale of evolution of the SMBHB orbit due to the emission of GWs is compared to the corresponding characteristic timescale of orbit evolution defined by interaction with the accretion disc (see Eq. 15). For a circular orbit and small (see, e.g. Landau & Lifshitz 1975, Chapter 13, Section 110), the separation between two components in a binary system changes with time according to
where is the power carried away by GWs:
@@
This variation in corresponds to the timescale
From , and taking into account Eq. 15 and 19, we find that the radiation of GWs plays a main role in the orbital evolution of SMBHB when
As follows from Eq. 14 and 20, we cannot exclude that the evolution of the source J02041415 is already determined by the radiation of GWs because . The corresponding timescale of orbit evolution, as follows from (14) and (19), is
@@
Since the statistical distributions of major physical parameters (masses and mass ratios, spins, all orbital parameters) of the SMBHBs as a class of objects are largely unknown, we refrain from making any other predictions in this paper about the GW manifestations of SMBHBs for shorter time intervals. Nevertheless, in the next two subsections we present some considerations on the possibility of detecting GWs from individual SMBHBs by laser interferometer antennas such as LISA (Colpi et al. 2023) (in subsection 4.4) and the possible contribution of SMBHBs to the GW background detected by PTAs (Agazie et al. 2023, and references therein) (in subsection 4.5). These considerations outline the contours of future studies of SMBHBs as multi-messenger objects.
4.4 Detection of gravitational waves generated by an individual SMBHB
In this subsection, we touch briefly on the future detection of GWs by LISA-type detectors and give some rough constraints on SMBHB parameters for which such detection is possible. Let us assume that a prospective GW detector can detect all GWs if (i) their frequencies are in the range
and (ii) their amplitudes, (such an amplitude is called the strain of a GW), are larger than some value, (called the sensitivity of the detector).
Such an assumption is a rough replacement for the non-monotonic dependence of the sensitivity of any real GW detector on the frequency of the receiving signal (see e.g. Colpi et al. (2023) for LISA and Agazie et al. (2023) for NANOgrav). The main purpose of such an assumption is not only to simplify the restrictions given below on the parameters of SMBHBs, but also to make the physical meaning of these restrictions clearer.
@@ 4.4 Detection of gravitational waves generated by an individual SMBHB
Another constraint comes from a comparison of the amplitude of the emitted GW with . To make an order-of-magnitude estimate, we can use the quadrupole formula (see Landau & Lifshitz 1975, Chapter 13, Section 110). Omitting tensor indices we get
where is the second time derivative of the quadrupole moment, is the orbital angular velocity, and is the luminosity distance from the SMBHB to the detector of GWs. Then one can obtain the following inequality:
@@
The last, very simple constraint on comes from the well-known fact that the radius of the last stable circular orbit around a Schwarzschild black hole is . Hence,
@@
Finally, we have the following constraints on the parameters of SMBHBs generating detectable GWs (see also Ivanov et al. 2015):
@@ 4.5 SMBHBs and GW background detected by pulsar timing arrays
This subsection is devoted to the assumed population of SMBHBs that contributes to the GW background radiation detected by PTAs (Agazie et al. 2023, and references therein). Let us assume that the population of SMBHBs is evolving with the characteristic time given by Eq. 15, and consists of objects with approximately the same parameters. The number of SMBHBs, , that could make a significant contribution to the GW background detected by PTAs above a given sensitivity threshold, , and in a given frequency range (Eq. 22) is , where . As follows from Eq. 15 and 19,
The SMBHB separation, , still satisfies inequalities 26 and 30, while inequality 29 should be significantly modified. Indeed, for background radiation, inequality 28 should be rewritten as
@@
The constraint on can now be obtained as a condition for the compatibility of inequalities 37 and 26:
Hence,
@@
It is notable that the parameter in our extremely simplified toy model does not appear in the above constraint for detectability of GWs by PTAs, contrary to the constraint in Eq. 33 for detectability of individual SMBHB GWs generated. The characteristic time of the existence of close-enough SMBHBs is proportional to (see Eq. 15), while the characteristic time when individual SMBHBs contribute to the background radiation is inversely proportional to (see Eq. 19). Hence, when we deal with the GW background, the fraction of contributing SMBHBs is proportional to (see Eq. 35). For this reason, the corresponding requirement for the amplitude of the GW, which itself is proportional to (see Eq. 29), is weakened by a factor of (see Eq. 35 and 36). The fact that the intensity of background GW radiation generated by SMBHBs does not depend on (or depends weakly in a model more accurate than the toy model presented in this paper) favours the formation of SMBHBs from less numerous but more massive SMBHs and more numerous but less massive SMBHs.
@@
5 Outlook for future direct imaging studies of SMBHBs as multi-messengers
@@
As was shown above, getting decisive information on the parameters of an SMBHB as a generator of GWs requires its measurable parameters to be obtained at the microarcsecond scale, at least indirectly. Arguably, the most convincing evidence and in-depth study opportunity would be provided by direct imaging observations of potential SMBHB objects with a microarcsecond resolution, in which the duality of the source’s morphology would be visible directly, as in several cases mentioned in Section 1 but at much smaller separations of milliparsecs – see the estimates presented in Section 4. This can be illustrated following a very convenient visual representation presented in Fig. 1 by D’Orazio et al. (2024): the dual AGN currently occupying the very top edge of the plot with the rest-frame periods of the order of yr should be ‘pushed down’ by five orders of magnitude to orbital periods of the order of years in the area of VLBI astrometry.
@@
Microarcecond angular resolution is beyond reach of any existing astronomical imaging instrument or facility in any domain of the electromagnetic spectrum. The currently record-holding Earth-based VLBI instruments cannot provide such a resolution due to the fundamental limitations of the longest available baselines, which cannot exceed the Earth’s diameter, and the Earth’s atmosphere opacity at wavelengths shorter than about 0.7 mm. The only plausible solution for achieving the required angular resolution is to place a VLBI system in space, above the Earth’s atmosphere and with no limitations on the baseline length. Several concepts of such systems have been developed over the past decade. One of them, THEZA (TeraHertz Exploration and Zooming-in for Astrophysics) has been presented in response to the European Space Agency’s call for inputs into the prospective strategic program Voyage 2050 (Gurvits et al. 2021, 2022). It suggests a multi-element space-borne VLBI configuration able to operate at millimetre and sub-millimetre wavelengths with an angular resolution of the order of 1 as. Another concept called BHEX (Black Hole Explorer) considers an Earth–Space VLBI configuration operating at millimetre/sub-millimetre wavelengths (Johnson et al. 2024). With a bit of luck, and accepting the prediction by D’Orazio et al. (2024, section 3.1), the currently ongoing analysis of the RadioAstron non-imaging AGN Survey (Kovalev et al. 2020), reaching in the best cases an angular resolution of , might offer one to several new SMBHB candidates. However, they will still require decisive imaging observations with future space-borne VLBI systems.
@@
Another potentially efficient multi-messenger approach to investigating the contribution of mergers of black hole binaries into the GW background at frequencies higher than the ones explored by PTA, accessible to LISA, is suggested by Ellis et al. (2024). Their model reconciles the pre-JWST and JWST observations of the co-growth of masses of stellar population of galaxies and their central black holes, including SMBHBs. The quantitative estimates of GW generation by evolving SMBHBs presented in section 4 might become useful in assessing the observational perspectives of LISA and other GW instruments covering specific frequency ranges.
@@
While discussing prospects of future detections of SMBHBs as ‘visually’ binary AGNs, we need to underline the difference between traditional morphological VLBI studies of jets in AGNs, especially their innermost bases, and the SMBHs themselves. The latter would become accessible for direct imaging observations if the technique demonstrated by the millimetre-wavelength VLBI Event Horizon Telescope (EHT) in the detection of the black hole shadows in M 87∗ (EHT 2019) and Sgr A∗ (EHT 2022) improved its angular resolution at least by an order of magnitude. Again, this would become possible only with prospective space-borne extensions of EHT such as BHEX and THEZA.